40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Tue Feb 22 07:53:06 2022, Yehonathan, Update, WOPO, Waking up WOPO 
    Reply  Fri Feb 25 17:09:53 2022, shruti, Update, WOPO, Waking up WOPO - green beam 
       Reply  Wed Mar 2 14:38:55 2022, Yehonathan, Update, WOPO, Waking up WOPO - some more fiddling and a plan 
          Reply  Fri Mar 4 11:49:45 2022, shruti, Update, WOPO, Waking up WOPO - attempts at readout LO_shot_noise.pdf
             Reply  Thu Mar 10 17:12:01 2022, awade, Update, WOPO, Waking up WOPO - attempts at readout 
                Reply  Wed Mar 16 12:22:44 2022, Yehonathan, Update, WOPO, Waking up WOPO - attempts at readout Moku_Noise.pdfPD_Connected_no_light.pdfDiff_Channel.pdf
                   Reply  Tue Mar 22 09:19:38 2022, shruti, Update, WOPO, Waking up WOPO - back to green path before_opt.jpgaround_waveguide.png
Message ID: 2728     Entry time: Fri Mar 4 11:49:45 2022     In reply to: 2727     Reply to this: 2731
Author: shruti 
Type: Update 
Category: WOPO 
Subject: Waking up WOPO - attempts at readout 

[Yehonathan, Shruti]

1. Doubling cavity and green beam

Since we had left the lasers ON with the shutters closed we wanted to see if the powers measured after opening the shutter would be similar to what it was when we left. We realized that opening and closing the green shutter destabilizes the doubling cavity (the FI is after the shutter and the shutter does not seem to be a good dump), which in turn changes the SHG crystal temperature (possibly because of the power fluctuation within the crystal). Re-opening the shutter requires some tuning of the temperature and offset to recover similar output power. Finally, after some tuning, we were able to see 156 mW of green light.

 

2. Attempt at measuring LO shot noise

  • We want to measure the BHD output A-B channel, which we expect to be dominated by shot noise since all the classical noise would be canceled when optimally balanced, but found that one of the PDs was inverted so that the sum of the two channels would be what we needed to measure. Since the SR 560 has only an A-B option, we used a second SR 560 to invert the B channel before subtracting
  • Operating at an LO power of ~4 mW did not give us sufficient clearance from the dark noise in each channel with the SR 560s, which was around the max power I believe we're supposed to use for the BHD LO
  • Somehow measuring the output directly, without the SR 560, gave us some clearance over the dark noise at ~1 MHz and higher (possibly because 1 MHz is the SR560's BW) so we decided to measure the time series of both channels together and do the optimized subtraction and FFT offline

3. Subtracted noise signal spectra [Attachment 1]

  • The plots show the noise spectra of the channels measured individually. The 'gain adjusted' means that A was multiplied by 1.05 and B by 0.95 in order to get the two plots to more or less line up
  • We used the Moku to measure the timeseries at a sampling rate of 10.2 MS/s for a period of 1.2 ms with AC coupling and 50 Ohm impedance. Elog 2324 suggests the designed measurement was for a 50 Ohm load so we should be impedance matched but I'm yet to convince myself
  • Our estimate for the shot noise was 77 nV/rtHz for 4 mW of power and 87% QE, using a TI gain of 2kOhm (the black dashed line in the plot). If we were impedance matched the yellow trace must be higher than this estimate
  • In our next measurements, we will also record the dark noise, carefully measure the power. There is obviously sonething wrong with the plot

 

 

30 Mar 22 edit: script here, data here

Attachment 1: LO_shot_noise.pdf  18 kB  Uploaded Fri Mar 4 17:13:42 2022  | Hide | Hide all
LO_shot_noise.pdf
ELOG V3.1.3-